# When to Truncate the Archive? On the Effect of the Truncation Frequency in Multi-Objective Optimisation

### v1

@@ Abstract.
Zhiji Cui, Zimin Liang and Miqing Li (corresponding author) are with the School of Computer Science, University of Birmingham, Edgbaston, Birmingham, UK (emails: zxc990@student.bham.ac.uk; zxl525@student.bham.ac.uk; m.li.8@bham.ac.uk).
Lie Meng Pang and Hisao Ishibuchi are with the Southern University of Science and Technology, Shenzhen, Guangdong, China. (emails: panglm@sustech.edu.cn; hisao@sustech.edu.cn)
@@ 1. Introduction
Conventionally, the archive truncation can be performed either once a new solution arrives or a batch of new solutions arrive. The former corresponds to the case of updating the population once a new solution is generated (i.e., the $\mu+1$ evolution mode) in MOEAs such as SMS-EMOA (Beume2007, ) and MOEA/D (Zhang2007, ), while the latter corresponds to the case of updating the population once a number of solutions are generated (i.e., the $\mu+\mu$ evolution mode) in MOEAs such as in NSGA-II (Deb2002, ), IBEA (Zitzler2004, ) and NSGA-III (Deb2014, ). In addition, recently people tend to use an unbounded archive to store all nondominated solutions found during the search (Fieldsend2003, ; Ishibuchi2020, ; chen4902239updating, ; Rudolph2016, ; bezerra2019archiver, ; brockhoff2019benchmarking, ; li2024empirical, ; pang2020algorithm, ; bian2024archive, ). Afterwards, the archive is truncated down to a small, but representative subset provided to the decision maker (as a large number of solutions can easily overwhelm them). Intuitively, one may think this approach is better since we have all the information in hand before performing the truncation. The $\mu+1$ approach that truncates the archive once every new solution arrives may be less promising due to the unknown about future input when the decision is made.
@@ 4.1. NSGA-II
Lastly, it is worth mentioning that regardless of the truncation approaches, uniformity of the solutions obtained by NSGA-II is poor. This is because the crowding distance fails to measure the crowdedness level of solutions in the space with three or more objectives – an issue which has been frequently reported and studied in the area (see, e.g., (Kukkonen2006, ; Li2014a, )).

### v2

@@ 1. Introduction
Conventionally, the archive truncation can be performed either once a new solution arrives or a batch of new solutions arrive. The former corresponds to the case of updating the population once a new solution is generated (i.e., the $\mu+1$ evolution mode) in MOEAs such as SMS-EMOA (Beume2007, ) and MOEA/D (Zhang2007, ), while the latter corresponds to the case of updating the population once a number of solutions are generated (i.e., the $\mu+\mu$ evolution mode) in MOEAs such as in NSGA-II (Deb2002, ), IBEA (Zitzler2004, ) and NSGA-III (Deb2014, ). In addition, recently people tend to use an unbounded archive to store all nondominated solutions found during the search (Fieldsend2003, ; Ishibuchi2020, ; chen4902239updating, ; Rudolph2016, ; bezerra2019archiver, ; brockhoff2019benchmarking, ; li2024empirical, ; pang2020algorithm, ; bian2024archive, ). Afterwards, the archive is truncated down to a small, but representative subset provided to the decision maker (as a large number of solutions can easily overwhelm them). Intuitively, one may think this approach is better since we have all the information in hand before performing the truncation. The $\mu+1$ approach that truncates the archive once every new solution arrives may be less promising due to the unknown about future input when the decision is made (Lopez2011, ).
@@ 4.1. NSGA-II
Lastly, it is worth mentioning that regardless of the truncation approaches, uniformity of the solutions obtained by NSGA-II is poor. This is because the crowding distance fails to measure the crowdedness level of solutions in the space with three or more objectives – an issue which has been frequently reported and studied in the area, theoretically (zheng_runtime_2024, ) and empirically (Kukkonen2006, ; Li2014a, ).